Steam with a flow rate of 1360 kg/h enters an adiabatic nozzle (adiabatic, Q=0) at 1378

kPa, specific volume of 0.147 m3


/kg, and specific internal energy of 2510 kJ/kg. The exit


conditions are 137.8 kPa, 1.099 m3


/kg specific volume, and 2263 kJ/kg of specific internal


energy. Find the exit velocity.


1
Expert's answer
2021-11-11T12:03:16-0500

Solution;

For an adiabatic nozzle;

W=0

Q=0

Given;

"\dot{m}=1360kg/h=0.3778kg/s"

"P_1=1378kPa"

"v=0.147m^3\/kg"

"u_1=2510kJ\/kg"

"P_2=137.8kPa"

"v_2=1.099m^3\/kg"

"u_2=2263kJ\/kg"

"V_1=3.05m\/s"

Energy balance equation;

"h_1+\frac{v_1^2}{2000}=h_2+\frac{v_2^2}{2000}"

But h=u+PV

"u_1+P_1v_1+\frac{V_1^2}{2000}=u_2+P_1v_1+\frac{V_2^2}{2000}"

By direct substitution;

"2510+1378×0.147+\frac{3.05^2}{2000}=2263+137.8×1.099+\frac{V_2^2}{2000}"

"V_2^2=621409.3025"

"V_2=788.3m\/s"